Write an english version of the following equations;
a) 2x-5=x+1
b) 2(x+2)=14
c) n+5=n/2-6
d) 7-3a = 5+a

Answer by tutordave(1)

a) 2x-5=x+1
First subtract x from both sides...
x-5=1
Then add 5 to both sides...
x=6 (ANSWER)
b) 2(x+2)=14
Divide both sides by 2...
x+2=7
Subtract 2 from both sides...
x=5 (ANSWER)
c) n+5=n/2-6
Multiply both sides by 2...
2(n+5)=n-12
Distribute the left-hand side...
2n+10=n-12
Subtract n from both sides...
n+10=-12
Subtract 10 from both sides...
n=-22 (ANSWER)
d) 7-3a = 5+a
Add 3a to both sides...
7 = 5+4a
Subtract 5 from both sides...
2 = 4a
Divide both sides by 4
2/4 = a
Switch sides, and simplify the fraction...
a = 1/2 (ANSWER)